Kayla has 25 ounces of a 75% boric acid solution. How many ounces of 35% boric acid solution should she add to obtain a solution of 60% boric acid?
Answer by stanbon(75887)

Kayla has 25 ounces of a 75% boric acid solution. How many ounces of 35% boric acid solution should she add to obtain a solution of 60% boric acid?
-------
Equation:
acid + acid = acid
0.75*25 + 0.35x = 0.60*(25+x)
-----
75*25 + 35x = 60*25 + 60x
=========
25x = 15*25
x = 15 ounces (amt. of 35% ba to be added)
======
Cheers,
Stan H.
========